OpenStudy (howard-wolowitz):

hard question:

OpenStudy (solomonzelman):

ok, tell me what function has a vertical asymptote x=0 (the y-axis) and horizontal asymptote y=0 (the x-axis) ?

OpenStudy (solomonzelman):

or, at least try to think of the SIMPLEST example of a function where \(x\ne0\).

OpenStudy (howard-wolowitz):

alright here: a rational function

OpenStudy (solomonzelman):

good, but give me a specific function pliz.

OpenStudy (howard-wolowitz):

ok

OpenStudy (howard-wolowitz):

1/x

OpenStudy (howard-wolowitz):



OpenStudy (solomonzelman):

yes, y=1/x is good

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm f}(x)=\frac{\rm 1}{x} }\)

OpenStudy (solomonzelman):

in this function \(x\ne0\), and thus, \(x=0\) is the vertical asymptote of this function.

OpenStudy (solomonzelman):

now, will this function ever have a y-value of 0?

OpenStudy (howard-wolowitz):

yes. I think because of the 1

OpenStudy (solomonzelman):

no, because 1/x is never going to be 0, for any value of x. (and if x is 0, then it's undefined)

OpenStudy (solomonzelman):

and since y is never equal to 0 therefore horizontal asymptote is \(y=0\)

OpenStudy (howard-wolowitz):

so then yes it will have a value of 0

OpenStudy (solomonzelman):

no, it won't have a a value of y=0, because 1/x is same as y and 1/x is never zero

OpenStudy (solomonzelman):

ok, y=1/x will be something like this